There are three environments for constructing lists. In each one each new item is begun with an \item command. In the itemize environment the start of each item is given a marker, in the enumerate environment each item is marked by a number. These environments can be nested within each other in which case the amount of indentation used is adjusted accordingly:

\begin{itemize}
\item Itemized lists are handy.
\item However, don't forget
  \begin{enumerate}
  \item The `item' command.
  \item The `end' command.
  \end{enumerate}
\end{itemize}

• Itemized lists are handy.

• However, don’t forget

1. The ‘item’ command.

2. The ‘end’ command.

The third list making environment is description. In a description you specify the item labels inside square brackets after the \item command. For example:

Three animals that you should
know about are:
\begin{description}
  \item[gnat] A small animal...
  \item[gnu] A large animal...
  \item[armadillo] A ...
\end{description}

Three animals that you should know about are:

gnat A small animal that causes no end of trouble.

gnu A large animal that causes no end of trouble.

armadillo A medium-sized animal.